The Mathclub, VIT-AP wants to conduct a group event for its members. So the club president has to fix the group size the event. When he tries to fix the size to be 5 members in each group, 4 members are left; when he tries to fix the size to be 6 members in each group, 5 members are left; When he fixes the size to be 7 members in each group, 6 members are left. What is the smallest number of members that the club has?


1
Expert's answer
2021-01-04T20:45:41-0500

Let "x" be the number of members that the club has. Then we have the following system of linear congruences:


"\begin{cases}
x\equiv 4\ (\mod 5)\\
x\equiv 5\ (\mod 6)\\
x\equiv 6\ (\mod 7)
\end{cases}"


The first congruence is equivalent to the equality "x=4+5t,\ t\in\mathbb Z." Put this in the second congruence:


"4+5t\equiv 5\ (\mod 6)"


"5t\equiv 1\ (\mod 6)"


"-t\equiv 1\ (\mod 6)"


"t\equiv -1\ (\mod 6)"


Consequently, "t=-1+6s,\ s\in\mathbb Z."


"x=4+5t=4+5(-1+6s)=-1+30s"


Put "x" in the third congruence of the system:


"-1+30s\equiv 6\ (\mod 7)"


"30s\equiv 7\ (\mod 7)"


"30s\equiv 0\ (\mod 7)"


"s\equiv 0\ (\mod 7)"


Therefore, "s=7k,\ k\in\mathbb Z."


We conclude that


"x=-1+30s=-1+30\cdot 7k=-1+210k, k\in\mathbb Z."


For we have the smallest positive "x=-1+210=209".


Therefore, the smallest number of members that the club has is 209.